OpenStudy (anonymous):

Urgent Help needed due today pls help In a hydrogen atom in its ground state, the electron is on average a distance of about 1×10-10 m from the proton. What is the magnitude of the electric field due to the proton a distance of 0.9 10-10 m away?

OpenStudy (john_es):

Using, \[|E|=\frac{1}{4\pi\epsilon_0}\frac{Q_{proton}}{r^2}=\frac{1}{4\pi\epsilon_0}\frac{Q^2_{electron}}{r^2}=\frac{1}{4\pi 8.85\cdot10^{-12}}\frac{1.609\cdot10^{-19}}{(0.9\cdot10^{-10})^2}=2.87\cdot10^{-8}\ N/C \]

OpenStudy (john_es):

Units: N/C